The first-order rate constant for the decomposition of N2O5 to NO2 and O2 at 70 °C is 5.3 × 10–3 s–1. Suppose we start with 0.500 mol of N2O5(g) in a 0.700 L container, how many moles of N2O5 will remain after 1.5 min?



1
Expert's answer
2021-04-05T07:58:00-0400



"ln(\frac{A_0}{A_t})=kt"


"ln(\frac{A_0}{A_t})=5.3×10^{-3}×1.5×60"


"ln(\frac{A_0}{A_t})=0.447"


"A_t=(\frac{0.500}{e^{0.447}})=0.31 moles"